OpenStudy (anonymous):

\[Calculate: \int\limits_{}^{}\sin^2x*dx,\] By rewriting the identity:\[\cos(2x)=1-2*\sin^2(x)\]

OpenStudy (callisto):

∫(sinx)^2 dx = ∫ (1-cos2x)/2 dx = (1/2) ∫ (1-cos2x) dx = (1/2) [ x - (1/2) sin2x] +c = (1/2)x - (1/4) sin2x +c

OpenStudy (lalaly):

\[\sin^2x=\frac{1}{2}(1-\cos(2x))\]so the integral becomes\[\int\limits{\frac{1}{2}(1-\cos(2x))}dx\]\[=\int\limits{\frac{1}{2}-\frac{1}{2}\cos(2x)}dx\]\[=\int\limits{\frac{1}{2}dx }-\int\limits{\frac{1}{2}\cos(2x)dx}=\frac{1}{2}x-\frac{1}{4}\sin(2x)+C\]

OpenStudy (anonymous):

Thanks! :)